OpenStudy (itrymath):

Given the function g(x) = -3x + 4, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values. (3 points) The value of g(-2) is smaller than the value of g(4). The value of g(-2) is the same as the value of g(4). The values of g(-2) and g(4) cannot be compared. The value of g(-2) is larger than the value of g(4).

OpenStudy (itrymath):

@mathmate

OpenStudy (itrymath):

@3mar

OpenStudy (3mar):

Well, I am here. What did you get?

OpenStudy (3mar):

It is just a substitution process!

OpenStudy (itrymath):

no its not

OpenStudy (itrymath):

i got D

OpenStudy (itrymath):

@3mar hello?

OpenStudy (3mar):

Yes I am here

OpenStudy (3mar):

\[\Large g(x) = -3x + 4\] \[\Large g(-2) = -3(-2) + 4=6+4=10\] \[\Large g(4) = -3(4) + 4=-12+4=-8\] You compare the two result and pick your correct answer!

OpenStudy (itrymath):

OOOOOHHH i forgot !!!

OpenStudy (itrymath):

omg

OpenStudy (itrymath):

i still think answer D is the correct answer

OpenStudy (3mar):

\[\Huge\color{Simon}\checkmark\]

OpenStudy (3mar):

You hit the right target!

OpenStudy (itrymath):

yayya

OpenStudy (itrymath):

thanks

OpenStudy (3mar):

You are welcome! Did you get what you are looking for?